OpenStudy (anonymous):

A cube has an edge length of 2x. Write an expression for the total area of the cube please help

OpenStudy (anonymous):

how to i get the expression

OpenStudy (ybarrap):

Total surface area is area of one face times 6 So Area = 6 * (2x)^2 = 6*4x^2 = 24x^2

OpenStudy (ybarrap):

For example, if the cube had length equal to 1, then 2x = 1, so x = 1/2. Then area of one face is 1*1=1, since there are 6 faces, total area should be 6, using our formula. So plugging in 1/2 into the equation, gives you 6, like it should.

OpenStudy (anonymous):

i thought you dont count the bottom face

OpenStudy (ybarrap):

There's no reason to exclude it.

OpenStudy (anonymous):

thanks